find the elimination method
6x-6y=-48
-3x-2y=-11

Answer by stanbon(75887)

find the elimination method
6x-6y=-48
-3x-2y=-11
------
Multiply thru the 2nd equation by 2 to get:
6x -6y = -48
-6x -4y = -22
-------------------------
Add and solve for "y":
-10y = -70
y = 7
----
Solve for "x":
6x - 6y = -48
6x -6*7 = -48
6x = -6
x = -1
============
Cheers,
Stan H.
================